OpenStudy (anonymous):

State the x- intercept and y- intercept for the graph of each equation. Then graph the equation using x- and y-intercepts 1) 2x + 3y = 12 2) y = -2x -4 I hope someone maybe able to help me so I understand I submitted these two above and I got them wrong so I hope someone maybe able to help me so I understand as I am not sure what I have done wrong. Thanks

OpenStudy (aripotta):

to find the x-intercept, substitute 0 for y in the equation and solve for x. to find the y-intercept, substitute 0 for x in the equation and solve for y. then to graph, just graph the intercepts on the coordinate plane, and then draw a line through the points

OpenStudy (anonymous):

you plug in the y equation into the x one. 2x+3(-2x-4)=12 2x+-6x-12=12 -4x-12=12 -4x=24 x=-6, then plug in -6 into the y equation. y= -2(-6)-4 y= 12-4 y= 8 you should have (-6,8)

OpenStudy (anonymous):

I am confused I got (-6, 4) and for number two I have (-2. -4) not sure what I did wrong

OpenStudy (aripotta):

x-intercept is going to be (x, 0) and y-intercept is going to be (0, y)

OpenStudy (aripotta):

do what i said, and you should be all set

OpenStudy (anonymous):

I think I understand a little better Thanks

OpenStudy (aripotta):

1) 2x + 3y = 12 to find x-intercept: 2x + 3(0) = 12 solve for x to find y-intercept: 2(0) + 3y = 12 solve for y

OpenStudy (anonymous):

Thanks